OpenStudy (anonymous):

long division 0.14 )¯¯¯645

OpenStudy (anonymous):

got a calculator?

OpenStudy (anonymous):

yes but I want to know how to do the work

OpenStudy (goldeverything):

645\[645\div0.14\]

OpenStudy (anonymous):

oh, then start with \[64500\div14\]

OpenStudy (anonymous):

do whatever long division is necessary to get the answer

OpenStudy (anonymous):

$$\frac{645}{\frac{14}{100}}=\frac{64,500}{14}=4607+\frac{1}{7}$$

OpenStudy (texaschic101):